• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name rsholmes

    Log entry time 16:27:44 on May 16, 2010

    Entry number 322830

    keyword=Run Coordinator report, May 15 Swing - May 16 Day

    Swing shift started with hall in access for left dipole. After Jack left, beam was restored and Moller crew took it for the last half of Swing and all of Owl.

    Day shift Jack spent working on the left dipole again, still unsuccessfully.

    We plan to run 1-arm production during tonight's Swing and Owl. Day shift Jack would like to get assistance from Accelerator power supply techs to attempt left dipole repair.



    A copy of this log entry has been emailed to: rom,kkumar@physics.mass.edu,souder@physics.syr.edu,rsholmes@physics.syr.edu,crowder